OpenStudy (anonymous):

Use the information to evaluate and compare Δy and dy. (Round your answers to four decimal places.) y = x^4 + 9 x = −3 Δx = dx = 0.01 Please Check my work, I think I'm doing this right... Δy= f(x+ Δx)-f(x) =f(-3+0.01)-f(-3) =f(-2.99)-f(-3) =4(-2.99)^3-4(-3)^3 =1.076404 dy=f^1(x)dx=f^1(-3)(0.01)=4(-.03)^3 =-1.08e^-4

OpenStudy (aum):

\[ \frac{dy}{dx} = 4x^3 \\ \left[ \frac{dy}{dx} \right]_{x=-3} = 4(-3)^3 = -108\\ dy \approx = -108 * dx = -108 * 0.01 = -1.08 \]

OpenStudy (anonymous):

thanks, i see where i went wrong

OpenStudy (aum):

you are welcome.

OpenStudy (aum):

\[ \Delta y = f(x+\Delta x) - f(x) = (-3+0.01)^4 + 9 - (-3)^4 - 9 = -1.0746 \]

OpenStudy (anonymous):

so for Δy, you don't use the derivative? you just use the original function?

OpenStudy (aum):

Yes. Delta y is the change in y for a small change delta x in x.

OpenStudy (aum):



OpenStudy (aum):

delta y = f(x+deltax) - f(x)